OpenStudy (anonymous):

Determine the equation of the line, in point-slope form, that runs through points (2, -4) and (-4, -6) A) y-4=1/3(x+2) B) y+4=1/3x-2 C) y+4=1/3(x-2)

OpenStudy (uri):

Find slope first \[m =\frac{ (y2-y1) }{ (x2-x1)}\]

OpenStudy (britanyashton):

plug in your known points, into the slope formula.

OpenStudy (britanyashton):

that is your first step.

OpenStudy (anonymous):

Slope is -2/-6

OpenStudy (britanyashton):

Great. Simplify that fraction now.

OpenStudy (uri):

formula is y-y1=mx(x-x1)

OpenStudy (britanyashton):

When you do you have a numerator and denomerator that are both positive.

OpenStudy (anonymous):

1/3

OpenStudy (uri):